Grade 12Physical Chemistry

How to calculate the no. of particles in hcp crystal lattice?

Jayant Motwani
9 Years agoGrade 12

1 Answer

cool matrix
9 Years ago
Total no. of particles in hcp lattice is 6. This is bcoz in a hcp structure there are three layers having two hexagonal sheets and a triangle between them. Of these three particles at the vertices of triangle of the middle sheet lie completely inside the lattice. Now the vertices of hexagon contain 6 particles but they r shared by another 6 lattice so contribution per lattice is 6*1/6 that is 1. but as i said earlier there r two hexahonal layers so 1*2= 2. and finally the centre of these two hexagonal layers also contain atoms that r shared by 2 lattices so contribution per lattice is 1/2 but as there r two hexagonal layrs so 1/2*2=1. So finally total atoms per hcp lattice will be (3+ 2+ 1) = 6. Hope u got else be free toask and kindly approve if u like it.